# Identification of Additive Continuous-time Systems in Open and Closed-loop

### v1

@@ 1 Introduction
The goal in system identification is to obtain mathematical descriptions of systems using input and output data. These models can play a critical role for prediction, analysis, and design of control laws for electrical, mechanical, biological, or environmental systems, among others [42, 27]. A distinction can be made between discrete-time and continuous-time system identification methods. Discrete-time identification algorithms provide models valid only at the sampling instants, while continuous-time identification methods use data to estimate models that represent the true system at any point in time.
@@
Linear continuous-time system identification methods [37] are widely used and successful in a range of practical applications [48, 13, 10], offering several advantages over the standard discrete-time algorithms. One advantage of continuous-time methods is their ability to directly incorporate a priori knowledge of the relative degree of the physical systems they model. This feature is particularly useful in estimating, e.g., mechanical systems, as they often exhibit no impulse response discontinuities due to the double integration relationship between force and position [14], leading to models with relative degree equal to 2. In contrast, discrete-time methods must also account for sampling zeros that are not present in the continuous-time system representation [2], requiring an additional optimization step if the resulting discrete-time model is later converted to continuous-time [24]. Recent research has proposed tailored variants of instrumental variables to address this particular challenge [22].
@@
When addressing the identification of physical systems, increasingly stringent performance requirements necessitate the use of parsimonious models, i.e., models with the fewest number of parameters that can adequately describe the phenomenon under investigation. From a statistical standpoint, it is well known that parsimonious models result in reduced variance of the estimated models compared to over-parameterized models [42]. Most linear continuous-time identification methods parameterize the model structure as an unfactored transfer function with a user-defined number of poles and zeros, which can limit the flexibility of the model. This is the case for methods such as the Simplified Refined Instrumental Variable method for Continuous-time systems (SRIVC, [50]), the Poisson moment functional approach [39], the Least-Squares State-Variable Filter method (LSSVF, [47]), and other linear filter and integral methods [11]. The closed-loop variants of these estimators are also limited to estimating unfactored transfer functions [15, 16, 49]. Although more general model structures can be incorporated using the prediction error method and maximum likelihood paradigm [1], typically one unfactored transfer function is estimated in the linear and time-invariant continuous-time identification problem [28].
@@
While unfactored transfer functions are standard choices for the model structure in linear system identification, many practical applications related to, e.g., flexible motion systems [29, 31] and vibration analysis [14], often involve systems that are more easily interpreted as a sum of transfer functions with distinct denominators, typically corresponding to different resonant modes. In addition, modal parameter estimation for structures subject to vibrations is essential for design and model validation, to guarrantee safety and quality control [46, 38]. Additive model parametrizations have advantages such as leading to physically more insightful models for fault diagnosis [6] and improving the numerical conditioning of parameter estimation for high-order or highly-resonant systems [19]. These model parametrizations, which have been considered in statistics [26] and econometrics [25, 32], offer increased model flexibility and the ability to decentralize the analysis of each additive component for optimization and control purposes. Despite some contributions in non-linear discrete-time finite-impulse response and generalized Hammerstein model estimation [3, 4], the application of additive model structures in the realm of system identification has been limited. Recently, [23] introduced a block coordinate algorithm to identify continuous-time systems under an additive model structure. However, this approach is confined to open-loop setups and can be computationally demanding due to the need for estimating each submodel with a tailored version of the SRIVC method at each iteration of the descent algorithm.
@@
The remainder of this paper is structured as follows. Section 2 introduces the problem setup for both open and closed-loop settings. Section 3 describes the optimality conditions for the additive model structures of both settings. Section 4 contains the proposed unified iterative procedure for additive continuous-time system identification, and its asymptotic analysis can be found in Section 5. Extensive simulations can be found in Section 6, while Section 7 contains experimental setup results. Concluding remarks are presented in Section 8. A technical lemma used for proving the main theoretical results can be found in the Appendix.
@@ 2 System and Model Setup
where $p$ is the Heaviside operator (i.e., $pu(t)=\frac{\textnormal{d}}{\textnormal{d}t}u(t)$) and $u(t)$ is the input signal. Each subsystem $G_{i}^{*}(p)$ has $n_{i}$ poles and $m_{i}$ zeros, and can be expressed as $B_{i}^{*}(p)/A_{i}^{*}(p)$, where the numerator and denominator polynomials are assumed coprime, i.e., they do not share roots. These polynomials are given by
@@
with $a_{n_{i}}^{*}\neq 0$ and $m_{i}\leq n_{i}$. We assume without loss of generality that the $A_{i}^{*}(p)$ polynomials are anti-monic, i.e., their constant coefficient is fixed to 1, and that they are jointly coprime. In addition, for obtaining a unique characterization of $\{G_{i}^{*}(p)\}_{i=1}^{K}$, we assume that at most one subsystem $G_{i}^{*}(p)$ is biproper. The polynomials $A_{i}^{*}(p)$ and $B_{i}^{*}(p)$ are jointly described by the parameter vector
@@
A noisy measurement of the output is retrieved at every time instant $t=t_{k}$, $k=1,\dots,N$, where $\{t_{k}\}_{k=1}^{N}$ are evenly spaced in time. That is,
@@
where $v(t_{k})$ is assumed to be a zero-mean stationary random process of finite variance $\sigma^{2}$. Two frameworks are considered in this work: open and closed-loop identification. Block diagrams that describe these setups are presented in Figure 1. The output noise $v(t_{k})$ is assumed to be uncorrelated with the input samples $u(t_{k})$ for the open-loop case, and uncorrelated with the reference signal $r(t_{k})$ for the closed-loop case. For both settings, we assume that the input signal $u(t)$ is constant between samples, i.e., it has a zero-order hold (ZOH) behavior. The closed-loop setting considers an input signal generated from a known discrete-time LTI controller $C_{\textnormal{d}}(q)$, with $q$ being the forward-shift operator. Formally, we can describe the input $u(t_{k})$ in terms of the reference and output noise as follows:
@@ 3.2 Closed-loop setting
In contrast to the open-loop analysis, an output-error loss function of the form in (7) applied to closed-loop system identification can result in an asymptotically biased estimator. This phenomenon is due to the possible correlation of the output noise with the input $u(t)$ through the closed-loop interconnection [45, 20]. The bias in closed-loop can be mitigated or reduced completely as $N$ tends to infinity if an instrumental variable approach is considered, as in [17] for the unfactored transfer function case. Thus, instead of (7), for the closed-loop setting we are interested in computing the instrumental variable solution
@@
where $\bm{\zeta}(t_{k})\in\mathbb{R}^{K+\sum_{i=1}^{K}(n_{i}+m_{i})}$ is an instrument vector that is assumed to be uncorrelated with the output noise $v(t_{k})$. A crucial aspect of the estimator in (11) is how the instrument vector is chosen for yielding consistent estimates of minimum covariance. Conditions on the instrument vector under which the estimator is generically consistent are provided in Lemma 3.1.
@@
Assume that the instrument vector $\bm{\zeta}(t_{k})$ is uncorrelated with the output noise $v(t_{k})$, and that $r(t_{k})$ and $v(t_{l})$ are stationary and mutually independent for all integers $k$ and $l$. Then, the estimator (11) is generically consistent if $\mathbb{E}\left\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\right\}$ is generically non-singular [^1] In this context, a statement $s$, which depends on the elements $\bm{\rho}$ of some open set $\Omega\subseteq\mathbb{R}^{n}$, is generically true with respect to $\Omega$ [41] if the set $M=\{\bm{\rho}\in\Omega|s(\bm{\rho})\textnormal{ is not true}\}$ has Lebesgue measure zero in $\Omega$. with respect to the system and model denominator parameters, where $\tilde{\bm{\varphi}}^{r}(t_{k})$ is formed by stacking the vectors $\{\tilde{\bm{\varphi}}_{i}^{r}(t_{k})\}_{i=1}^{K}$, where
@@
Proof. As the sample size tends to infinity, the sum in $k$ in (11) converges to an expected value due to the stationarity assumptions on the reference and noise signals [40], which leads to $\bar{\bm{\theta}}_{i}=\lim_{N\to\infty}\hat{\bm{\theta}}_{i}$ satisfying
$\mathbb{E}\left\{\bm{\zeta}(t_{k})\left(y(t_{k})-\sum_{i=1}^{K}G_{i}(p,\bar{\bm{\theta}}_{i})u(t_{k})\right)\right\}$ $=\mathbf{0}$
$\iff\sum_{i=1}^{K}\mathbb{E}\big{\{}\bm{\zeta}(t_{k})\left[G_{i}^{*}(p)-G_{i}(p,\bar{\bm{\theta}}_{i})\right]u(t_{k})\big{\}}$ $=\mathbf{0}$
$\iff\sum_{i=1}^{K}\mathbb{E}\big{\{}\bm{\zeta}(t_{k})\left[G_{i}^{*}(p)-G_{i}(p,\bar{\bm{\theta}}_{i})\right]\tilde{r}(t_{k})\big{\}}$ $=\mathbf{0},$ (13)
@@
where $\tilde{\bm{\varphi}}_{i}^{r}(t_{k})$ is given in (12), and $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$ is a Sylvester matrix that is non-singular for $i=1,\dots,K$ due to the coprimeness of the polynomials $A_{i}^{*}(p)$ and $B_{i}^{*}(p)$ [41, Lemma A3.1]. If the vectors $\tilde{\bm{\varphi}}^{r}(t_{k})$ and $\bm{\eta}$ are formed by stacking the vectors $\tilde{\bm{\varphi}}_{i}^{r}(t_{k})$ and $\bm{\eta}_{i}$ respectively, then (13) reduces to the condition
$\mathbb{E}\left\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\right\}\begin{bmatrix}\mathbf{S}(-B_{1}^{*},A_{1}^{*})&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\mathbf{S}(-B_{K}^{*},A_{K}^{*})\\ \end{bmatrix}^{-\top}\bm{\eta}=\mathbf{0},$
which implies that $\bm{\eta}$ is generically the zero vector (i.e., $G_{i}(p,\bar{\bm{\theta}}_{i})=G_{i}^{*}(p)$ for all $i=1,\dots,K$) if the matrix $\mathbb{E}\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\}$ is generically non-singular. $\square$
@@
In accordance to Lemma 3.1, we will focus solely on instrument vectors that meet the non-singularity requirement, which defines a family of generically consistent estimators. Among these, we aim to characterize the one that obtains the least asymptotic covariance. In cases where the transfer function is unfactored, it is established in both discrete-time [44, 43] and continuous-time [18] formulations that the instrument vector achieving the lower bound of the asymptotic covariance matrix for the parameter estimator is given by the noise-free component of the regressor vector. However, for the additive model description, the computation of $\hat{\bm{\beta}}$ cannot be carried out in a straightforward analytical manner, and the model output cannot be expressed using just one regressor vector. Despite these difficulties, Theorem 3.2 confirms that the asymptotic distribution of $\hat{\bm{\beta}}$ in Eq. (11) follows a normal distribution, and the asymptotic covariance can be explicitly computed.
@@
Consider the estimator (11). Assume that $r(t_{k})$ and $v(t_{l})$ are stationary and mutually independent for all integers $k$ and $l$, and that the instrument vector $\bm{\zeta}(t_{k})$ is constructed such that $\hat{\bm{\beta}}$ is a consistent estimator of $\bm{\beta}^{*}$. Then, $\hat{\bm{\beta}}$ is asymptotically Gaussian distributed, i.e.,
@@
$\mathbf{P}_{\textnormal{IV}}$ $=\sigma^{2}\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}$
×$\times\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\mathbb{E}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1},$
@@
#$\#\hskip 1.42271pt\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\bm{\zeta}(t_{k})v(t_{k}).$ (16)
@@
$=\sigma^{2}\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}.$
@@
$\mathbb{E}\left\{\begin{bmatrix}\bm{\zeta}(t_{k})\\ \bm{\psi}(t_{k})\end{bmatrix}\begin{bmatrix}\bm{\zeta}(t_{k})\\ \bm{\psi}(t_{k})\end{bmatrix}^{\top}\right\}\succeq\mathbf{0}.$ (18)
The non-singularity $\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}$ implies that the matrix $\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}$ is positive definite. Thus, (18) is equivalent to the Schur complement
$\begin{split}&\mathbb{E}\left\{\bm{\psi}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}-\mathbb{E}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\\ &\hskip 2.84544pt\times\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1}\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}\succeq\mathbf{0}.\end{split}$ (19)
Provided that the matrices above are non-singular, (19) is equivalent to
$\mathbb{E}\left\{\bm{\psi}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}\preceq\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}$
×$\hskip 19.91684pt\times\mathbb{E}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\mathbb{E}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1}.$
Equality holds when $\bm{\zeta}(t_{k})=\mathbf{M}\bm{\psi}(t_{k})$, with $\mathbf{M}$ being a non-singular constant matrix. Since $\mathbf{M}$ can be canceled out from (11), we set $\mathbf{M}$ equal to the identity matrix without loss of generality, which leads to the desired result. $\square$
Corollary 3.3 indicates that the optimal instrument vector is given by a filtered version of the reference signal that depends on the true system parameters, as seen in (15). Since the true parameters are not known, the solution for the unfactored transfer function case adopted in, e.g., the CLSRIVC estimator [16], is to let $\bm{\zeta}(t_{k})$ depend on $\bm{\beta}$ as well. In the additive identification case, we let the instrument vector take the form (9), but where, for $i=1,\dots,K$,
@@
[…]

### v2

# Identification of Additive Continuous-time Systems in Open and Closed Loop

@@ Abstract
[…]
Rodrigo A. González et al., Identification of Additive Continuous-time Systems in Open and Closed Loop
To appear in Automatica, uploaded to Arxiv on September 25th, 2024
@@ 1 Introduction
The goal in system identification is to obtain mathematical descriptions of systems using input and output data. These models can play a critical role for prediction, analysis, and design of control laws for electrical, mechanical, biological, or environmental systems, among others [45, 31]. A distinction can be made between discrete-time and continuous-time system identification methods using sampled data. Identification algorithms for discrete-time systems provide difference equation models that give meaningful information only at the sampling instants. In contrast, continuous-time system identification methods use sampled data to estimate differential equation models that represent the system at any point in time.
@@
Linear continuous-time system identification methods [40] are widely used and successful in a range of practical applications [52, 14, 11], offering several advantages over the standard discrete-time algorithms. One advantage of continuous-time methods is their ability to directly incorporate a priori knowledge of the relative degree of the physical systems they model. This feature is particularly useful in estimating, e.g., mechanical systems, as they often exhibit no impulse response discontinuities due to the double integration relationship between force and position [15], leading to models with relative degree equal to 2. In contrast, discrete-time methods must also account for sampling zeros that are not present in the continuous-time system representation [2], requiring an additional optimization step if the resulting discrete-time model is later converted to continuous time [27]. Recent research has proposed tailored variants of instrumental variables to address this particular challenge [25].
@@
When addressing the identification of physical systems, increasingly stringent performance requirements necessitate the use of parsimonious models, i.e., models with the fewest number of parameters that can adequately describe the phenomenon under investigation. From a statistical standpoint, it is well known that parsimonious model structures that contain the true system within the model set result in reduced variance compared to over-parameterized model sets [47]. Most linear continuous-time identification methods parameterize the model structure as an unfactored transfer function, that is, a quotient of polynomials in the Laplace transform variable $s$ with numerator and denominator polynomial degrees indicating the number of zeros and poles of the model, respectively. Unfactored transfer function parametrizations may not always offer the most parsimonious model structure, but due to their simplicity they are widely adopted in the Simplified Refined Instrumental Variable method for Continuous-time systems (SRIVC, [54]), the Poisson moment functional approach [42], the Least-Squares State-Variable Filter method (LSSVF, [51]), and other linear filter and integral methods [12]. The closed-loop variants of these estimators are also limited to estimating unfactored transfer functions [16, 17, 53]. Although more general model structures can be incorporated using the prediction error method and maximum likelihood paradigm [1], typically one unfactored transfer function is estimated in the linear and time-invariant continuous-time identification problem [32].
@@
While unfactored transfer functions are standard choices for the model structure in linear system identification, many practical applications related to, e.g., flexible motion systems [35] and vibration analysis [15], often involve systems that are more easily interpreted as a sum of transfer functions with distinct denominators, typically corresponding to different resonant modes. In addition, modal parameter estimation for structures subject to vibrations is essential for design and model validation, to guarrantee safety and quality control [50, 41]. Additive model parametrizations have advantages such as leading to physically more insightful models for fault diagnosis [6, 7] and improving the numerical conditioning of parameter estimation for high-order or highly-resonant systems [20]. These model parametrizations, which have previously been considered in statistics [29] and econometrics [28], offer increased model flexibility and the ability to decentralize the analysis of each additive component for optimization and control purposes. Despite some contributions in nonlinear discrete-time finite-impulse response and generalized Hammerstein model estimation [3, 4], the application of additive model structures in the realm of system identification has been limited. Recently, [26] introduced a block coordinate algorithm to identify continuous-time systems under an additive model structure. However, this approach is confined to open-loop setups and can be computationally demanding due to the need for estimating each submodel with a tailored version of the SRIVC method at each iteration of the descent algorithm [21].
@@
The remainder of this paper is structured as follows. Section 2 introduces the problem setup for both open and closed-loop settings. Section 3 describes the optimality conditions for the additive model structures of both settings. Section 4 contains the proposed unified iterative procedure for additive continuous-time system identification with its asymptotic analysis. Extensive simulations can be found in Section 5, while Section 6 contains experimental setup results. Concluding remarks are presented in Section 7. A technical lemma used for proving the main theoretical results can be found in the Appendix.
@@ 2 System and Model Setup
where $p$ is the Heaviside or derivative operator (i.e., $pu(t)=\frac{\textnormal{d}}{\textnormal{d}t}u(t)$) and $u(t)$ is the input signal. Each subsystem $G_{i}^{*}(p)$ has $n_{i}$ poles and $m_{i}$ zeros, and can be expressed as $B_{i}^{*}(p)/A_{i}^{*}(p)$, where the numerator and denominator polynomials are assumed coprime, i.e., they do not share roots. These polynomials are given by
@@
with $a_{n_{i}}^{*}\neq 0$ and $m_{i}\leq n_{i}$. We assume, without loss of generality, that the $A_{i}^{*}(p)$ polynomials do not share roots and that they are anti-monic, i.e., their constant coefficient is fixed to 1. In addition, to obtain a unique characterization of $\{G_{i}^{*}(p)\}_{i=1}^{K}$, we assume that at most one subsystem $G_{i}^{*}(p)$ has the same number of poles and zeros. The polynomials $A_{i}^{*}(p)$ and $B_{i}^{*}(p)$ are jointly described by the parameter vector
@@
A noisy measurement of the output is retrieved at every time instant $t=t_{k}$, $k=1,\dots,N$, where $\{t_{k}\}_{k=1}^{N}$ are evenly spaced in time [^1] Extensions of our proposed identification methods for irregularly sampled signals are feasible by following similar steps to the ones found in [30].. That is,
@@
where $v(t_{k})$ is assumed to be a zero-mean stationary random process of finite variance $\sigma^{2}$. Two frameworks are considered in this work: open and closed-loop identification. Block diagrams that describe these setups are presented in Figure 1. The output noise $v(t_{k})$ is assumed to be uncorrelated with the input samples $u(t_{k})$ for the open-loop case, and uncorrelated with the reference signal $r(t_{k})$ for the closed-loop case. For both settings, we assume for simplicity that the input signal $u(t)$ is known to be constant between samples, i.e., it has a zero-order hold (ZOH) behavior. Other known intersample behaviors, such as first-order hold or band-limited, can also be addressed within the proposed framework at the expense of a more involved notation and ad-hoc prefiltering techniques [23]. The closed-loop setting considers an input signal generated from a known discrete-time LTI controller $C_{\textnormal{d}}(q)$, with $q$ being the forward-shift operator. Formally, we can describe the input $u(t_{k})$ in terms of the reference and output noise as follows:
@@
Remark 2.
Throughout this paper we assume that the input (for the open-loop case) and reference signals (for the closed-loop case) are quasi-stationary. Consequently, we will use the standard definition of expectation for quasi-stationary signals [31, p. 34]
$\overline{\mathbb{E}}\{g(t_{k})\}:=\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\mathbb{E}\{g(t_{k})\}.$
@@ 3.2 Closed-loop setting
In contrast to the open-loop analysis, an output-error loss function of the form in (7) applied to closed-loop system identification can result in an asymptotically biased estimator. This phenomenon is due to the correlation of the output noise with the input $u(t)$ through the closed-loop interconnection [49, 22]. The bias in closed loop can be mitigated or reduced completely as $N$ tends to infinity if an instrumental variable approach is considered, as in [18] for the unfactored transfer function case. Thus, instead of (7), for the closed-loop setting we are interested in computing the instrumental variable solution
@@
where $\bm{\zeta}(t_{k})\in\mathbb{R}^{K+\sum_{i=1}^{K}(n_{i}+m_{i})}$ is an instrument vector that is assumed to be uncorrelated with the output noise $v(t_{k})$. Such assumption is not restrictive in practice, since the instrument vector is typically formed by filtered versions of the reference signal, which are designed by the user.
A crucial aspect of the estimator in (11) is how the instrument vector is chosen for yielding consistent estimates of minimum covariance. Conditions on the instrument vector under which the estimator is generically consistent are provided in Lemma 3.1.
@@
Assume that the instrument vector $\bm{\zeta}(t_{k})$ is uncorrelated with the output noise $v(t_{k})$, and that $r(t_{k})$ and $v(t_{l})$ are quasi-stationary and mutually independent for all integers $k$ and $l$. Then, the estimator (11) is generically consistent if $\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\right\}$ is generically nonsingular [^2] In this context, a statement $s$, which depends on the elements $\bm{\rho}$ of some open set $\Omega\subseteq\mathbb{R}^{n}$, is generically true with respect to $\Omega$ [44] if the set $M=\{\bm{\rho}\in\Omega|s(\bm{\rho})\textnormal{ is not true}\}$ has Lebesgue measure zero in $\Omega$. with respect to the system and model denominator parameters, where $\tilde{\bm{\varphi}}^{r}(t_{k})$ is formed by stacking the vectors $\{\tilde{\bm{\varphi}}_{i}^{r}(t_{k})\}_{i=1}^{K}$, where
@@
Proof. As the sample size tends to infinity, the sum in $k$ in (11) converges to an expected value due to the quasi-stationarity assumptions on the reference and noise signals [43], which leads to $\bar{\bm{\theta}}_{i}=\lim_{N\to\infty}\hat{\bm{\theta}}_{i}$ satisfying
$\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\left(y(t_{k})-\sum_{i=1}^{K}G_{i}(p,\bar{\bm{\theta}}_{i})u(t_{k})\right)\right\}$ $=\mathbf{0}$
$\iff\sum_{i=1}^{K}\overline{\mathbb{E}}\big{\{}\bm{\zeta}(t_{k})\left[G_{i}^{*}(p)-G_{i}(p,\bar{\bm{\theta}}_{i})\right]u(t_{k})\big{\}}$ $=\mathbf{0}$
$\iff\sum_{i=1}^{K}\overline{\mathbb{E}}\big{\{}\bm{\zeta}(t_{k})\left[G_{i}^{*}(p)-G_{i}(p,\bar{\bm{\theta}}_{i})\right]\tilde{r}(t_{k})\big{\}}$ $=\mathbf{0},$ (13)
@@
where $\tilde{\bm{\varphi}}_{i}^{r}(t_{k})$ is given in (12), and $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$ is a Sylvester matrix that is nonsingular for $i=1,\dots,K$ due to the coprimeness of the polynomials $A_{i}^{*}(p)$ and $B_{i}^{*}(p)$ [44, Lemma A3.1]. If the vectors $\tilde{\bm{\varphi}}^{r}(t_{k})$ and $\bm{\eta}$ are formed by stacking the vectors $\tilde{\bm{\varphi}}_{i}^{r}(t_{k})$ and $\bm{\eta}_{i}$ respectively, then (13) reduces to the condition
$\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\right\}\begin{bmatrix}\mathbf{S}(-B_{1}^{*},A_{1}^{*})&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\mathbf{S}(-B_{K}^{*},A_{K}^{*})\\ \end{bmatrix}^{-\top}\bm{\eta}=\mathbf{0},$
which implies that $\bm{\eta}$ is generically the zero vector (i.e., $G_{i}(p,\bar{\bm{\theta}}_{i})=G_{i}^{*}(p)$ for all $i=1,\dots,K$) if the matrix $\overline{\mathbb{E}}\{\bm{\zeta}(t_{k})\tilde{\bm{\varphi}}^{r\top}(t_{k})\}$ is generically nonsingular. $\square$
@@
In accordance to Lemma 3.1, we will focus solely on instrument vectors that meet the nonsingularity requirement, which defines a family of generically consistent estimators. Among these, we aim to characterize the one that obtains the least asymptotic covariance. In cases where the transfer function is unfactored, it is established in both discrete-time [48, 46] and continuous-time [19] formulations that the instrument vector achieving the lower bound of the asymptotic covariance matrix for the parameter estimator is given by the noise-free component of the regressor vector. However, for the additive model description, the computation of $\hat{\bm{\beta}}$ cannot be carried out in a straightforward analytical manner, and the model output cannot be expressed using just one regressor vector. Despite these difficulties, Theorem 3.2 confirms that the asymptotic distribution of $\hat{\bm{\beta}}$ in Eq. (11) follows a normal distribution, and the asymptotic covariance can be explicitly computed.
@@
Consider the estimator (11). Assume that $r(t_{k})$ and $v(t_{l})$ are quasi-stationary and mutually independent for all integers $k$ and $l$, and that the instrument vector $\bm{\zeta}(t_{k})$ is constructed such that $\hat{\bm{\beta}}$ is a consistent estimator of $\bm{\beta}^{*}$. Then, $\hat{\bm{\beta}}$ is asymptotically Gaussian distributed, i.e.,
@@
$\mathbf{P}_{\textnormal{IV}}$ $=\sigma^{2}\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}$
×$\times\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\overline{\mathbb{E}}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1},$
@@
#$\#\hskip 1.42271pt\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\bm{\zeta}(t_{k})v(t_{k}).$ (16)
@@
$=\sigma^{2}\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}.$
@@
$\overline{\mathbb{E}}\left\{\begin{bmatrix}\bm{\zeta}(t_{k})\\ \bm{\psi}(t_{k})\end{bmatrix}\begin{bmatrix}\bm{\zeta}(t_{k})\\ \bm{\psi}(t_{k})\end{bmatrix}^{\top}\right\}\succeq\mathbf{0}.$ (18)
The nonsingularity of $\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}$ implies that the matrix $\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}$ is positive definite. Thus, (18) is equivalent to the Schur complement
$\begin{split}&\overline{\mathbb{E}}\left\{\bm{\psi}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}-\overline{\mathbb{E}}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\\ &\hskip 2.84544pt\times\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1}\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}\succeq\mathbf{0}.\end{split}$ (19)
Provided that the matrices above are nonsingular, (19) is equivalent to
$\overline{\mathbb{E}}\left\{\bm{\psi}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}\preceq\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\psi}^{\top}(t_{k})\right\}^{-1}$
×$\hskip 19.91684pt\times\overline{\mathbb{E}}\left\{\bm{\zeta}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}\overline{\mathbb{E}}\left\{\bm{\psi}(t_{k})\bm{\zeta}^{\top}(t_{k})\right\}^{-1}.$
Equality holds when $\bm{\zeta}(t_{k})=\mathbf{M}\bm{\psi}(t_{k})$, with $\mathbf{M}$ being a nonsingular constant matrix. Since $\mathbf{M}$ can be canceled out from (11), we set $\mathbf{M}$ equal to the identity matrix without loss of generality, which leads to the desired result. $\square$
Corollary 3.3 indicates that the optimal instrument vector is given by a filtered version of the reference signal that depends on the true system parameters, as seen in (15). Since the true parameters are not known, the solution for the unfactored transfer function case adopted in, e.g., the closed-loop SRIVC estimator (CLSRIVC, [17]), is to let $\bm{\zeta}(t_{k})$ depend on $\bm{\beta}$ as well. In the additive identification case, we let the instrument vector take the form (9), but where, for $i=1,\dots,K$,
@@
To unify the notations pertaining to the open and closed loop cases, from now on we denote the instrument vector composed by stacking the vectors (20) for $i=1,\dots,K$ as $\hat{\bm{\varphi}}(t_{k},\bm{\beta})$. The difference between this vector and the one described by (10) will be clear from the context.